# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v2

@@ 3 Problem Formulation
We make the following tripartite assumption on the neural network function: (i) For any fixed $\mathbf{x}$, $\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz and is definable under some o-minimal structure that includes polynomials and exponential. [^2] We note that definability in some o-minimal structure is a mild technical assumption and is satisfied by all modern deep neural networks. It allows us to ensure convergence of certain gradient flow trajectories (Ji & Telgarsky, 2020). (ii) For all $c>0$, $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),$ for some $L>2$ (i.e., the network is $L$-homogenous). (iii) The gradient $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz.
@@ 4.1 Main Result
Let $\mathbf{w}_{0}$ be a fixed unit vector. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exists $T_{\epsilon},\tilde{B}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and solution $\mathbf{w}(t)$ of eq. 2 with initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, we have
@@ 4.1.1 Proof Sketch of Theorem 1
Theorem 1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We briefly describe the challenges in extending Theorem 1 to ReLU networks in Appendix D.
@@ 5.1 Leaky ReLU
If $\overline{{\mathbf{W}}}_{1:L}$ is a KKT point of eq. 10, then $\|{\mathbf{a}}_{l}\|_{2}^{2}=1/\sqrt{L}$, for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/L$.
  • Necessary and sufficient condition. Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
  • Necessary and sufficient condition. Suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $q_{l}{\mathbf{a}}_{l}={\mathbf{b}}_{l+1},$ for all $l\in[L-2]$, and $q_{L-1}{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, where $q_{l}\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
In the above theorem, $\{{\mathbf{a}}_{i},{\mathbf{b}}_{i}\}_{i=1}^{L-1}$ are vectors used to build the rank-one hidden weights. Our first result provides the norm of these vectors. Next, for $\alpha\neq 1$ and assuming the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, we present necessary and sufficient conditions, which can be divided in two parts. First is the alignment: ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$, which implies that the column-space of $\overline{{\mathbf{W}}}_{l}$ and row-space of $\overline{{\mathbf{W}}}_{l+1}$ must be aligned. The second part requires ${\mathbf{b}}_{1}$ to be a KKT point of eq. 11, a smaller optimization problem, where note that $\sigma^{L-1}(\cdot)$ denotes the function resulting from composing $\sigma(\cdot)$ with itself $(L-1)$ times. The results for $\alpha=1$ are mostly similar, but do not require any non-negativity assumption.
@@ 5.2 Polynomial Leaky ReLU
If $\overline{{\mathbf{W}}}_{1:L}$ is a KKT point of eq. 10, then $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p}$, for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$.
  • Necessary and sufficient condition. Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{b}}_{l}={\mathbf{a}}_{l-1}/p^{1/4}$, for $12\leq l\leq L-$ 1, and $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, for some $q\in\{-1,1\}$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
  • Necessary and sufficient condition. Suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{b}}_{l}=q_{l}{\mathbf{a}}_{l-1}/p^{1/4},\overline{{\mathbf{w}}}=q_{L}{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, if $p$ is odd, and ${\mathbf{b}}_{l}=q_{l}|{\mathbf{a}}_{l-1}|/p^{1/4},\overline{{\mathbf{w}}}=q_{L}|{\mathbf{a}}_{L-1}|/(p\hat{p})^{1/4}$, if $p$ is even, where $q_{l}\in\{-1,1\}$, for all $2\leq l\leq L$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical absolute values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@ 5.3 Numerical Experiments
We conclude this section by again noting that under the rank-one and non-negativity assumption considered here, the sufficient conditions of Theorem 3 and 4 are also necessary conditions, and thus, the KKT points must be of the form described in those results. Since the rank-one and non-negativity assumption is satisfied for all the instances considered, it implies the KKT points are of form described in Theorem 3 and 4 for all the instances considered. Also, note that the constrained NCF is a non-convex problem and could have KKT points that don’t satisfy these assumptions, but gradient ascent seems to avoid them. Understanding this phenomenon is an interesting future direction, which would require analyzing the dynamics of gradient ascent.
@@ C.3 Proof of Theorem 4
Next, as shown in eq. 115, we have
${\mathbf{Z}}_{L}\propto({\mathbf{a}}_{L-1}^{p})^{\top}.$ (127)
Since $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$ and the entries of ${\mathbf{a}}_{L-1}$ are identical and non-negative, ${\mathbf{a}}_{L-1}^{p}\propto{\mathbf{a}}_{L-1}\propto\overline{{\mathbf{w}}}.$ Hence, ${\mathbf{Z}}_{L}\propto\overline{{\mathbf{W}}}_{L}.$
@@ Appendix D Challenges for non-differentiable neural networks
Recall that Theorem 1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We next describe the difficulty in extending Theorem 1 to ReLU networks. Consider a neural network that satisfies Assumption 1, except that the gradient is not assumed to be locally Lipschitz (this will include ReLU networks). Then, assuming square loss, we can define the gradient flow dynamics using the Clarke subdifferential as follows

### v3

@@ 3 Problem Formulation
We make the following tripartite assumption on the neural network function: (i) For any fixed $\mathbf{x}$, $\mathcal{H}(\mathbf{x};\mathbf{w})$ is locally Lipschitz and is definable under some o-minimal structure that includes polynomials and exponential. [^2] We note that definability in some o-minimal structure is a mild technical assumption and is satisfied by all modern deep neural networks. It allows us to ensure convergence of certain gradient flow trajectories (Ji & Telgarsky, 2020). (ii) For all $c>0$, $\mathcal{H}(\mathbf{x};c\mathbf{w})=c^{L}\mathcal{H}(\mathbf{x};\mathbf{w}),$ for some $L>2$ (i.e., the network is $L$-homogenous). (iii) $\mathcal{H}(\mathbf{x};\mathbf{w})$ is differentiable in ${\mathbf{w}}$ and its gradient, $\nabla\mathcal{H}(\mathbf{x};\mathbf{w})$, is locally Lipschitz.
@@ 4.1 Main Result
Let $\mathbf{w}_{0}$ be a fixed unit vector. Under Assumption 1, for any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exists $T_{\epsilon},\tilde{B}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and solution $\mathbf{w}(t)$ of eq. 2 with initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, we have
@@ 4.1.1 Proof Sketch of Theorem 1
Theorem 1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We briefly describe the challenges in extending Theorem 1 to ReLU networks in Appendix D, which also includes an empirical evidence that early directional convergence also occurs in ReLU networks.
@@ 5.1 Leaky ReLU
  1. (i)
    Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $\|{\mathbf{a}}_{l}\|_{2}^{2}=1/\sqrt{L}$, for all $l\in[L-1],$ $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/L$, ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
  2. (ii)
    Next, suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $\|{\mathbf{a}}_{l}\|_{2}^{2}=1/\sqrt{L}$, for all $l\in[L-1],$ $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/L$, $q_{l}{\mathbf{a}}_{l}={\mathbf{b}}_{l+1},$ for all $l\in[L-2]$, and $q_{L-1}{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, where $q_{l}\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
In the above theorem, $\{{\mathbf{a}}_{i},{\mathbf{b}}_{i}\}_{i=1}^{L-1}$ are vectors used to build the rank-one hidden weights. In both cases, we first provide the norm of these vectors. Next, for $\alpha\neq 1$ and assuming the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, we present additional necessary and sufficient conditions, which can be divided in two parts. First is the alignment: ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$, which implies that the column-space of $\overline{{\mathbf{W}}}_{l}$ and row-space of $\overline{{\mathbf{W}}}_{l+1}$ must be aligned. The second part requires ${\mathbf{b}}_{1}$ to be a KKT point of eq. 11, a smaller optimization problem, where note that $\sigma^{L-1}(\cdot)$ denotes the function resulting from composing $\sigma(\cdot)$ with itself $(L-1)$ times. The results for $\alpha=1$ are mostly similar, but do not require any non-negativity assumption.
@@ 5.2 Polynomial Leaky ReLU
  1. (i)
    Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p}$, for all $l\in[L-1],$ $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$, ${\mathbf{b}}_{l}={\mathbf{a}}_{l-1}/p^{1/4}$, for $12\leq l\leq L-$ 1, and $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, for some $q\in\{-1,1\}$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@
  2. (ii)
    Suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p}$, for all $l\in[L-1],$ $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$, ${\mathbf{b}}_{l}=q_{l}{\mathbf{a}}_{l-1}/p^{1/4},\overline{{\mathbf{w}}}=q_{L}{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, if $p$ is odd, and ${\mathbf{b}}_{l}=q_{l}|{\mathbf{a}}_{l-1}|/p^{1/4},\overline{{\mathbf{w}}}=q_{L}|{\mathbf{a}}_{L-1}|/(p\hat{p})^{1/4}$, if $p$ is even, where $q_{l}\in\{-1,1\}$, for all $2\leq l\leq L$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical absolute values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
@@ 5.3 Numerical Experiments
We conclude this section by again noting that under the rank-one and non-negativity assumption considered here, the sufficient conditions of Theorem 3 and 4 are also necessary conditions, and thus, the KKT points must be of the form described in those results. Since the rank-one and non-negativity assumption is satisfied for all the instances considered, it implies the KKT points are of form described in Theorem 3 and 4 for all the instances considered. Also, note that the constrained NCF is a non-convex problem and could have KKT points that do not satisfy these assumptions, but gradient ascent seems to avoid them. Understanding this phenomenon is an interesting future direction, which would require analyzing the dynamics of gradient ascent.
@@ 6 Conclusion and Future Directions
Acknowledgments
The authors graciously acknowledge gift funding from InterDigital, which partially supported this work.
@@ C.3 Proof of Theorem 4
Next, as shown in eq. 115, we have ${\mathbf{Z}}_{L}\propto({\mathbf{a}}_{L-1}^{p})^{\top}.$ Since $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$ and the entries of ${\mathbf{a}}_{L-1}$ are identical and non-negative, ${\mathbf{a}}_{L-1}^{p}\propto{\mathbf{a}}_{L-1}\propto\overline{{\mathbf{w}}}.$ Hence, ${\mathbf{Z}}_{L}\propto\overline{{\mathbf{W}}}_{L}.$
@@ Appendix D Challenges for non-differentiable neural networks
Recall that Theorem 1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We first provide an empirical evidence that early directional convergence also occurs in ReLU networks, and then briefly describe the challenges in extending Theorem 1 to ReLU networks.
Let $\sigma(x)=\max(x,0)$. We train the following $3-$homogeneous ReLU network $\mathcal{H}(\mathbf{x};\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{v})=\mathbf{v}^{\top}\sigma(\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x})))$, where $\mathbf{W}_{1}\in\mathbb{R}^{20\times 20}$, $\mathbf{W}_{2}\in\mathbb{R}^{30\times 20}$ and $\mathbf{v}\in\mathbb{R}^{30}$. For training, we minimize the square loss with respect to the output of a smaller neural network $\mathcal{H}^{*}(\mathbf{x};\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*},\mathbf{v}_{*})=10\cdot\mathbf{v}_{*}^{\top}\sigma(|\mathbf{W}_{2}^{*}|\sigma(\mathbf{W}_{1}^{*}\mathbf{x}))$, where the entries of $\mathbf{W}_{1}^{*}\in\mathbb{R}^{2\times 20}$, $\mathbf{W}_{2}^{*}\in\mathbb{R}^{2\times 2},\mathbf{v}_{*}\in\mathbb{R}^{2\times 1}$ are drawn from standard normal distribution. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{20}$. We define $\mathbf{w}$ to be the vector that contains all the entries of $\mathbf{W}_{1},\mathbf{W}_{2}\mbox{ and }\mathbf{v}$, $\widetilde{\mathbf{w}}=\mathbf{w}/\|\mathbf{w}\|_{2}$, and $\mathcal{N}(\mathbf{w})$ denotes the NCF for this case. The network is trained for 64900 iterations using gradient descent with step-size $5\cdot 10^{-3}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.01$ and $\mathbf{w}_{0}$ is a random unit norm vector. From Figure 4, we observe that the training loss does not change much, and the norm of the weights increases by a multiplicative factor but remains small. Also, from Figure 4, it is clear that the weights approximately converge in direction to a KKT point of the constrained NCF. In Figure 5, we plot the normalized absolute value of the weights at initialization and at iteration 64900. Similar to the experiment in Section 4.1.2, low-rank structure emerges among the hidden weights of neural networks. In fact, the two largest singular values of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}\mbox{ and }\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ are $(0.5741,0.0346)$ and $(0.5772,0.0067)$ respectively, indicating that they have approximately rank one.
Figure 4: Early training dynamics of a three layer ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))^{\top}\widetilde{\mathbf{w}}(t)/\|\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). The loss has barely changed, and the norm of weights remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
(a) At initialization
(b) At iteration 64900
Figure 5: Normalized absolute value of the weights at different stages of training
We next describe the difficulty in extending Theorem 1 to ReLU networks. Consider a neural network that satisfies Assumption 1, except that the gradient is not assumed to be locally Lipschitz (this will include ReLU networks). Then, assuming square loss, we can define the gradient flow dynamics using the Clarke subdifferential as follows